OpenStudy (anonymous):

An apple pie has a 12-inch diameter and the angle of one slice is 60 degrees. What is the area of that slice of pie? Use 3.14 for π

Directrix (directrix):

I think that will be: 60 is to the circle in degrees as x (slice area) is to area of circle.

Directrix (directrix):

60/360 = x/(pi*r^2) 1/6 = x/ (36* pi)

Directrix (directrix):

Cross multiply and solve for x. x = 6 pi Check that @Beechgrovestudent

Directrix (directrix):

Multiply 6 times 3.14 per the instructions.